Asked 04.06.2018

Gaseous methane ch4 will react with gaseous oxygen o2 to produce gaseous carbon dioxide co2 and gaseous water h2o . suppose 1.12 g of methane is mixed with 1.7g of oxygen. calculate the maximum mass of water that could be produced by the chemical reaction. round your answer to 3 significant digits
Answer #1

Molar mass of CH4,

MM = 1*MM(C) + 4*MM(H)

= 1*12.01 + 4*1.008

= 16.042 g/mol

mass(CH4)= 1.12 g

use:

number of mol of CH4,

n = mass of CH4/molar mass of CH4

=(1.12 g)/(16.04 g/mol)

= 6.982*10^-2 mol

Molar mass of O2 = 32 g/mol

mass(O2)= 1.7 g

use:

number of mol of O2,

n = mass of O2/molar mass of O2

=(1.7 g)/(32 g/mol)

= 5.312*10^-2 mol

Balanced chemical equation is:

CH4 + 2 O2 ---> 2 H2O + CO2

1 mol of CH4 reacts with 2 mol of O2

for 6.982*10^-2 mol of CH4, 0.1396 mol of O2 is required

But we have 5.312*10^-2 mol of O2

so, O2 is limiting reagent

we will use O2 in further calculation

Molar mass of H2O,

MM = 2*MM(H) + 1*MM(O)

= 2*1.008 + 1*16.0

= 18.016 g/mol

According to balanced equation

mol of H2O formed = (2/2)* moles of O2

= (2/2)*5.312*10^-2

= 5.312*10^-2 mol

use:

mass of H2O = number of mol * molar mass

= 5.312*10^-2*18.02

= 0.9571 g

Answer: 0.957 g